# SOFT GRAPH TRANSFORMER FOR MIMO DETECTION

### v3

@@ 3.3 Attention as Contextual Encoding and Message Passing
3.4 Soft-Input-Soft-Output Interface
In MIMO detection, the receiver often has access to soft priors from the decoder or other auxiliary sources. To incorporate this information in a principled way, we define a Soft-Input-Soft-Output (SISO) interface that operates on the dual-subgraph token representation introduced in Section 3.2.
Soft Input Embedding Block: This block processes two distinct types of soft-valued inputs:
  • Symbolic tokens: $\mathcal{T}_{\text{sym}}$ with dimensions $2N_{t},N_{\text{bits}}/2]$
  • Linear-constraint tokens: $\mathcal{T}_{\text{lin}}$ with dimensions $2N_{r},2N_{t}+2]$
Both input types are independently processed through dedicated Feed-Forward Network (FFN) embeddings:
  • Symbol embeddings: $\{\hat{x}_{\text{emb}}\}$ with dimensions $2N_{t},d_{\text{model}}]$
  • Channel observation embeddings: $\{{\tau}_{\text{emb}}\}$ with dimensions $2N_{r},d_{\text{model}}]$
Positional encoding is incorporated into the embedding process before the combined representation is forwarded to subsequent transformer layers.
Soft Output Block: This block transforms the processed embeddings back into soft outputs:
  • Receives embedded representations: $\{\hat{x}_{\text{emb}}\}$ with dimensions $2N_{t},d_{\text{model}}]$
  • Processes through FFN followed by Sigmoid activation
  • Produces final soft outputs: $\{\hat{x}_{i}\}_{i=1}^{2N_{t}}$ with original dimensions $2N_{t},N_{\text{bits}}/2]$
The architecture enables end-to-end soft information processing, where soft inputs are embedded, contextually transformed, and finally decoded into refined soft outputs suitable for iterative processing. The SISO interface provides flexibility to incorporate SGT with other signal denoising modules, such as SISO decoders or other detectors.
Table 1: Comparison of MIMO Detection Algorithms: Complexity, Structure, and Performance, Transformer-based MIMO has a cubic complexity trend because of its QR-decomposition pre-processing, OAMP/OAMPNet have a cubic complexity trend because of their matrix inverse operation.
Method Complexity Trend Performance
ML Detection $O(M^{N_{t}}),M=2^{N_{\text{bits}}}$ Exponential Upper bound
OAMP / OAMPNet $O(KN_{r}N_{t}^{2})$ Cubic Near-ML for large system
Transformer-based MIMO $O(N_{r}N_{t}^{2}+LN_{t}^{2}d_{\text{model}})$ Cubic Near-ML for small system
Proposed SGT (Ours) $L\cdot O(N_{r}^{2}+N_{t}^{2}+N_{r}N_{t})\cdot d_{\text{model}}$ Quadratic Better than Transformer-based MIMO for small system
Table 2: Runtime on a GPU with 24GB memory and 16384 parallel computing cores per 1000 samples (in seconds)
@@ 3.4 Soft-Input-Soft-Output Interface
4 Experiment
4.1 Ablation Studies
We assess the effectiveness of graph-aware tokenization and cross-attention in standalone MIMO detection under Perfect-CSI Rayleigh fading channels with Quadrature Phase-Shift Keying (QPSK) modulation. All models have a model dimension of 128 and an 8-layer architecture trained with the same learning rates, batch sizes, and training steps. The compared variants include:
  • Baseline: Encoder-only Transformer with QR preprocessing [9].
Graph-Aware Tokenization: Better Information Preservation
Cross-Attention: Faster and Better Convergence
@@ 4.2 Complexity Analysis
Small-Scale MIMO System
For small settings, runtime results in Table 2 show that SGT maintains runtime comparable to OAMPNet and Transformer-based MIMO, while delivering superior BER performance (Fig. 3). This suggests that SGT achieves the best trade-off between complexity and performance.
Large-scale MIMO System
We analyze the online inference complexity of representative MIMO detection algorithms in Tab. 1. Let $d_{\text{model}}$ be the Transformer embedding dimension, $N_{\text{bit}}$ the number of bits per symbol, and $L$ the number of encoder layers. In ultra-large-scale systems, SGT exhibits a more favorable quadratic complexity growth, making it more scalable than existing deep learning baselines. However, the detection performance under extremely large MIMO settings warrants further investigation.

### v4

@@ 3.3 Attention as Contextual Encoding and Message Passing
3.4 Soft-Input–Soft-Output Interface
In practical receivers, MIMO detectors often operate with soft priors provided by channel decoders or other auxiliary modules. To support iterative and modular receiver design, SGT is equipped with a soft-input–soft-output (SISO) interface that processes and outputs probabilistic information in a unified framework.
Specifically, SGT accepts two types of soft-valued inputs:
  • Prior symbol beliefs as symbolic tokens $\mathcal{T}_{\text{sym}}$;
  • Channel observations as linear-constraint tokens $\mathcal{T}_{\text{lin}}$.
These inputs are independently embedded and mapped into a shared latent space, where positional encodings are applied before being processed by the stacked Transformer layers.
At the output, the refined symbolic embeddings are projected back to bit-level posterior likelihoods through a lightweight feed-forward network, producing soft outputs compatible with downstream decoders. This design enables end-to-end soft information processing, where priors are iteratively refined through graph-aware attention mechanisms.
By explicitly supporting soft inputs and outputs, SGT can be seamlessly integrated with SISO decoders or other detection modules, making it a flexible building block for iterative receiver architectures.
4 Experiment
4.1 Ablation Studies
We assess the effectiveness of graph-aware tokenization and cross-attention in standalone MIMO detection under Perfect-CSI Rayleigh fading channels with Quadrature Phase-Shift Keying (QPSK) modulation. Training and model settings is shown in Tab. 1. The compared methods include:
  • Baseline: Encoder-only Transformer with QR preprocessing [1].
Table 1: Settings for Ablation Studies
Parameter Value
Number of SGT layers ($L$) 8
Number of attention heads 8
Model dimension ($d_{\text{model}}$) 128
Feed-forward 2-layer, ReLU, 128 hidden units.
Dropout rate 0.1
Constellation type QPSK
Training SNR range [5, 15][-5,\,15] dB
Number of training epochs 250,000
Optimizer AdamW
Learning rate $1\times 10^{-4}$
Training batch size 500 ($16\times 16$), 2500(else)
Graph-Aware Tokenization: Better Information Preservation
Cross-Attention: Faster and Better Convergence
Table 2: Runtime on GPU with 24GB memory and 16384 computing cores per 1000 samples (in seconds)
@@ 4.1 Ablation Studies
Table 3: Comparison of MIMO Detection Algorithms in Terms of Computational Complexity. Transformer-based MIMO exhibits cubic complexity due to QR-decomposition pre-processing, while OAMP/OAMPNet also shows cubic complexity due to matrix inversion operations.
Method Complexity
ML Detection $O(M^{N_{t}}),\;M=2^{N_{\text{bits}}}$
OAMP / OAMPNet $O(KN_{r}N_{t}^{2})$
Transformer-based MIMO $O(N_{r}N_{t}^{2}+LN_{t}^{2}d_{\text{model}})$
Proposed SGT (Ours) $L\cdot O(N_{r}^{2}+N_{t}^{2}+N_{r}N_{t})\cdot d_{\text{model}}$
@@ 4.2 Complexity Analysis
Small-Scale MIMO System: For small settings, runtime results in Table 2 show that SGT maintains runtime comparable to OAMPNet and Transformer-based MIMO, while delivering superior BER performance (Fig. 3). This suggests that SGT achieves the best trade-off between complexity and performance.
Large-scale MIMO System: We analyze the online inference complexity of representative MIMO detection algorithms in Tab. 3. Let $d_{\text{model}}$ be the Transformer embedding dimension, $N_{\text{bit}}$ the number of bits per symbol, and $L$ the number of encoder layers. In ultra-large-scale systems, SGT exhibits a more favorable quadratic complexity growth. However, the detection performance under extremely large MIMO settings warrants further investigation.